ELECTRONIC SPECTROSCOPY OF ZrCH IN THE VISIBLE REGION

\underline{A. J. MERER} AND J. R. D. PEERS, Chemistry Department, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; S. J. RIXON, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1.

Zirconium methylidyne, ZrCH(D) has been produced by reaction of CH_{4} or CD_{4} with laser-ablated zirconium in a free-jet expansion. The visible-region laser-induced fluorescence spectrum has been recorded. Wavelength-resolved fluorescence spectra have given the ZrCH(D) ground state frequencies ν_{2} (bend) = 580 (466) cm^{−1}and ν_{3} (Zr-C stretch) = 870 (810) cm^{−1}. Several perpendicular transitions are observed between 590 and 640 nm. Analysis of rotationally-resolved spectra of ZrCH and ZrCD will be presented and the geometry of these molecules will be discussed.